# Uncertainty Modeling in Graph Neural Networks via Stochastic Differential Equations

### v3

@@ 4.1 Out of Distribution Detection
Table 1 presents the AUROC and AURC scores for OOD detection across multiple datasets. AUROC evaluates the model’s ability to differentiate between in-distribution and out-of-distribution (OOD) samples, with higher scores indicating better discrimination. AURC measures the risk of misclassification as coverage increases, where lower values are preferred. The LGNSDE model (ours) consistently achieves the best AUROC and AURC scores across most datasets, indicating its superior performance in accurately identifying OOD samples and minimizing the risk of misclassification.
@@ 5 Conclusions and Future Work
In conclusion, LGNSDEs outperform the tested models, opening a new avenue for uncertainty quantification in graph data. In future work, stronger benchmarks should be included for a more comprehensive evaluation. Additionally, Neural SDEs face challenges with time and memory complexity. Further work should explore more scalable sampling methods to address these limitations.
@@ A.1 Notation
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ denote a graph with node set $\mathcal{V}$ and edge set $\mathcal{E}$. The node feature matrix at time $t$ is $\mathbf{H}(t)\in\mathbb{R}^{n\times d}$, where $n$ is the number of nodes and $d$ is the feature dimension. The evolution of $\mathbf{H}(t)$ is described by a Graph Neural Stochastic Differential Equation, with drift function $\mathbf{F}_{\mathcal{G}}(\mathbf{H}(t),t,\boldsymbol{\theta})$ and diffusion function $\mathbf{G}_{\mathcal{G}}(\mathbf{H}(t),t)$. Here, $\mathbf{F}_{\mathcal{G}}$ depends on the graph $\mathcal{G}$, the node features $\mathbf{H}(t)$, time $t$, and parameters $\boldsymbol{\theta}$. The diffusion function $\mathbf{G}_{\mathcal{G}}$ depends on $\mathcal{G}$ and $\mathbf{H}(t)$ but not on $\boldsymbol{\theta}$, as in practice, this is usually a constant function. The randomness is introduced through the Brownian motion $\mathbf{W}(t)$.
The constants $L_{f}$ and $L_{g}$ are Lipschitz constants for the drift and diffusion functions, respectively, ensuring the existence and uniqueness of the solution to the GNSDE. The linear growth condition is controlled by a constant $K$, preventing unbounded growth in $\mathbf{F}_{\mathcal{G}}$ and $\mathbf{G}_{\mathcal{G}}$. Finally, $\text{Var}(\mathbf{H}(t))$ represents the variance of the node features, capturing the aleatoric uncertainty in the system, which is also reflected in the variance of the model output $\mathbf{y}(t)=\mathbf{h}(\mathbf{H}(t))$.
@@ A.2 Technical Assumptions
$\|\mathbf{F}_{\mathcal{G}}(\mathbf{H}_{1}(t),t,\boldsymbol{\theta})-\mathbf{F}_{\mathcal{G}}(\mathbf{H}_{2}(t),t,\boldsymbol{\theta})\|_{F}$ $\leq L_{f}\|\mathbf{H}_{1}(t)-\mathbf{H}_{2}(t)\|_{F}$ (2)
@@
$\|\mathbf{F}_{\mathcal{G}}(\mathbf{H}(t),t,\boldsymbol{\theta})\|_{F}^{2}+\|\mathbf{G}_{\mathcal{G}}(\mathbf{H}(t),t)\|_{F}^{2}\leq K(1+\|\mathbf{H}(t)\|_{F}^{2}),$
@@ A.3 Proofs
Lemma 3.
Under assumptions 1-3, there exists a unique strong solution to a Graph Neural SDE of the form:
$d\mathbf{H}(t)=\mathbf{F}_{\mathcal{G}}(\mathbf{H}(t),t,\boldsymbol{\theta})\,dt+\mathbf{G}_{\mathcal{G}}(\mathbf{H}(t),t)\,d\mathbf{W}(t),$
whose variance bounds the variance of the model output $\mathbf{y}(t)$ as:
$\text{Var}(\mathbf{y}(t))\leq L^{2}_{h}\text{Var}(\mathbf{H}(t)),$
hence providing a meaningful measure of the total uncertainty.
@@
Using standard results from the theory of SDEs Pham [2009] applied to graph SDEs, it follows that the Lipschitz conditions of $\mathbf{F}_{\mathcal{G}}$ and $\mathbf{G}_{\mathcal{G}}$ ensure the existence and uniqueness of a strong solution $\mathbf{H}(t)$ to the GNSDE.
@@
Finally, assuming that the model output $\mathbf{y}(t)$ is a function of the latent state $\mathbf{H}(t)$, $\mathbf{y}(t)=\mathbf{h}(\mathbf{H}(t))$, where $\mathbf{h}:\mathbb{R}^{n\times d}\rightarrow\mathbb{R}^{n\times p}$ is a smooth function, we can apply Itô’s Lemma as follows:
$dy(t)=h^{\prime}(\mathbf{H}(t))\left[\mathbf{F}_{\mathcal{G}}(\mathbf{H}(t),t,\boldsymbol{\theta})\,dt+\mathbf{G}_{\mathcal{G}}(\mathbf{H}(t),t)\,d\mathbf{W}(t)\right]+\frac{1}{2}h^{\prime\prime}(\mathbf{H}(t))\mathbf{G}_{\mathcal{G}}(\mathbf{H}(t),t)^{2}\,dt.$
For the variance of $\mathbf{y}(t)$, we focus on the term involving $\mathbf{G}_{\mathcal{G}}(\mathbf{H}(t),t)\,d\mathbf{W}(t)$:
@@
Lemma 4.
Under assumptions 1-3, consider two initial conditions $\mathbf{H}_{0}$ and $\tilde{\mathbf{H}}_{0}=\mathbf{H}_{0}+\delta\mathbf{H}(0)$, where $\delta\mathbf{H}(0)\in\mathbb{R}^{n\times d}$ represents a small perturbation in the initial node features with $\|\delta\mathbf{H}(0)\|_{F}=\epsilon$. Assume that $\mathbf{H}_{0}$ is taken from a compact set $\mathcal{H}\subseteq\mathbb{R}^{n\times d}$. Then, the deviation between the solutions $\mathbf{H}(t)$ and $\tilde{\mathbf{H}}(t)$ of the Graph Neural SDE with these initial conditions remains bounded across time $t$, specifically:
@@
$d(\delta\mathbf{H}(t))=\left[\mathbf{F}_{\mathcal{G}}(\mathbf{H}_{1}(t),t,\boldsymbol{\theta})-\mathbf{F}_{\mathcal{G}}(\mathbf{H}_{2}(t),t,\boldsymbol{\theta})\right]dt+\left[\mathbf{G}_{\mathcal{G}}(\mathbf{H}_{1}(t),t)-\mathbf{G}_{\mathcal{G}}(\mathbf{H}_{2}(t),t)\right]d\mathbf{W}(t).$
@@
$d(\text{tr}(\delta\mathbf{H}(t)^{\top}\delta\mathbf{H}(t)))$ $=2\text{tr}\left(\delta\mathbf{H}(t)^{\top}\left[\mathbf{F}_{\mathcal{G}}(\mathbf{H}_{1}(t),t,\boldsymbol{\theta})-\mathbf{F}_{\mathcal{G}}(\mathbf{H}_{2}(t),t,\boldsymbol{\theta})\right]\right)dt$
@@
$\mathbb{E}[d(\text{tr}(\delta\mathbf{H}(t)^{\top}\delta\mathbf{H}(t)))]$ $=\mathbb{E}\left[2\text{tr}(\delta\mathbf{H}(t)^{\top}[\mathbf{F}_{\mathcal{G}}(\mathbf{H}_{1}(t),t,\boldsymbol{\theta})-\mathbf{F}_{\mathcal{G}}(\mathbf{H}_{2}(t),t,\boldsymbol{\theta})])\right]\,dt$
@@
A.4 GNSDE as a Continuous Representation of Graph ResNet with Stochastic Noise Insertion
Consider a Graph Neural Stochastic Differential Equation (GNSDE) represented as:

### v4

@@ 5.2 Out of Distribution Detection
Table 2 presents the AUROC and AURC scores for OOD detection across multiple datasets. AUROC evaluates the model’s ability to differentiate between in-distribution and out-of-distribution (OOD) samples, with higher scores indicating better discrimination. AURC measures the risk of misclassification as coverage increases, where lower values are preferred. LGNSDE consistently achieves the best AUROC and AURC scores across most datasets, indicating its superior performance in accurately identifying OOD samples and minimizing the risk of misclassification.
The accuracy was determined by applying an entropy-based threshold. Predictions with entropy above this threshold were classified as out-of-distribution (OOD), while those below were considered in-distribution. The optimal threshold was identified using a validation dataset, where it was selected to maximize overall classification performance.
5.3 Noise Perturbation
Metric Model Cora Citeseer Computers Photo Pubmed
AUROC (↑) GCN 0.9610 ± 0.0045 0.9096 ± 0.0056 0.9682 ± 0.0029 0.9909 ± 0.0002 0.8466 ± 0.0214
GNODE 0.9649 ± 0.0054 0.9077 ± 0.0062 0.9593 ± 0.0033 0.9892 ± 0.0007 0.8727 ± 0.0183
BGCN 0.9606 ± 0.0034 0.9069 ± 0.0076 0.9547 ± 0.0095 0.9845 ± 0.0028 0.7643 ± 0.0771
ENSEMBLE 0.9581 ± 0.0051 0.9166 ± 0.0048 0.9701 ± 0.0011 0.9893 ± 0.0004 0.8093 ± 0.0511
GRAPH GP 0.8979 ± 0.0050 0.8889 ± 0.0037 OOM OOM OOM
LGNSDE (Our) 0.9634 ± 0.0065 0.9172 ± 0.0070 0.9698 ± 0.0015 0.9911 ± 0.0004 0.8636 ± 0.0310
AURC (↓) GCN 0.0782 ± 0.0076 0.1755 ± 0.0079 0.0661 ± 0.0099 0.0185 ± 0.0005 0.2098 ± 0.0397
GNODE 0.0625 ± 0.0109 0.1676 ± 0.0085 0.1006 ± 0.0114 0.0209 ± 0.0018 0.1654 ± 0.0260
BGCN 0.0799 ± 0.0075 0.1687 ± 0.0122 0.1400 ± 0.0438 0.0359 ± 0.0059 0.3026 ± 0.0928
ENSEMBLE 0.0862 ± 0.0067 0.1690 ± 0.0099 0.0718 ± 0.0037 0.0230 ± 0.0008 0.2603 ± 0.0528
GRAPH GP 0.1827 ± 0.0081 0.2294 ± 0.0088 OOM OOM OOM
LGNSDE (Our) 0.0731 ± 0.0128 0.1612 ± 0.0145 0.0642 ± 0.0059 0.0184 ± 0.0010 0.1908 ± 0.0519
Accuracy (↑) GCN 0.8054 ± 0.0112 0.7162 ± 0.0145 0.8169 ± 0.0026 0.9139 ± 0.0013 0.6874 ± 0.0346
GNODE 0.8255 ± 0.0134 0.7213 ± 0.0116 0.7907 ± 0.0126 0.9074 ± 0.0061 0.7402 ± 0.0286
BGCN 0.7946 ± 0.0115 0.7034 ± 0.0223 0.7401 ± 0.0472 0.8754 ± 0.0180 0.5848 ± 0.0973
ENSEMBLE 0.7916 ± 0.0156 0.7199 ± 0.0139 0.8160 ± 0.0038 0.9091 ± 0.0013 0.6390 ± 0.0577
GRAPH GP 0.6486 ± 0.0124 0.6697 ± 0.0070 OOM OOM OOM
LGNSDE (Our) 0.8101 ± 0.0179 0.7214 ± 0.0178 0.8263 ± 0.0098 0.9165 ± 0.0022 0.7100 ± 0.0422
Table 3: AUROC (Mean ± Std), AURC (Mean ± Std), and Accuracy (Mean ± Std) for all datasets with noise perturbations. Red denotes the best-performing model, and blue denotes the second-best-performing model.
We evaluate the models with noise added during testing to assess their robustness to input perturbations. No noise is introduced during training or validation. At test time, Gaussian noise is applied to the input feature vectors. Specifically, the noisy inputs are defined as $\mathbf{X}_{\text{new\_test}}=\mathbf{X}_{\text{test}}+0.5\cdot\mathcal{N}(0,\sigma),$ where $\mathcal{N}(0,\sigma)$ represents element-wise Gaussian noise with mean 0 and standard deviation $\sigma$, independently applied to each input feature. Here, $\sigma$ is computed as the standard deviation of $\mathbf{X}_{\text{test}}$, i.e., $\sigma=\text{std}(\mathbf{X}_{\text{test}})$. By adding noise scaled by 0.5 times the standard deviation of the test set, we ensure that the perturbations are proportional to the feature distribution across all datasets.
Table 3 presents the results under the noisy perturbation setting, where our model, LGNSDE, consistently ranks among the top two across all datasets and metrics (AUROC, AURC, and Accuracy). This demonstrates its robust performance under noise, frequently outperforming other models. These results also align with the guarantees provided by Proposition 2, which predict bounded deviations under input perturbations, supporting the observed robustness of LGNSDE.
5.4 Active Learning
Figure 3: Active learning on the Cora dataset using two acquisition functions. The left plot shows a random selection of labels, while the right plot shows a selection based on maximum predictive entropy
We delve into decision-making under uncertainty in the context of an active learning setup, where the model selects its own training data. The experiments begin with the same set of observed labels as in the previous experiments (see Table 8). In each active learning round, 5 additional labels/nodes are incrementally included using an acquisition function, selected from either the validation or test dataset. After each new label is added, the models are trained for 25 epochs, and this process continues until the number of newly added labels has doubled.
Figure 3 shows the active learning experiment conducted on the Cora dataset nodes using two acquisition functions. In the right plot, labels are selected based on the highest predictive entropy, while in the left plot, they are selected randomly.
5.4.1 Random Acquisition Function
Metric Model Cora Citeseer Computers Photo Pubmed
AUROC (↑) GCN 0.9814 ± 0.0019 0.9447 ± 0.0017 0.9823 ± 0.0017 0.9933 ± 0.0006 0.9313 ± 0.0043
GNODE 0.9732 ± 0.0030 0.9149 ± 0.0080 0.9830 ± 0.0028 0.9923 ± 0.0008 0.9108 ± 0.0166
BGCN 0.9813 ± 0.0024 0.8339 ± 0.0341 0.9777 ± 0.0028 0.9896 ± 0.0023 0.7359 ± 0.0406
ENSEMBLE 0.9712 ± 0.0016 0.9139 ± 0.0047 0.9830 ± 0.0013 0.9926 ± 0.0008 0.9243 ± 0.0054
GRAPH GP 0.9434 ± 0.0050 0.9060 ± 0.0040 OOM OOM OOM
LGNSDE (Ours) 0.9812 ± 0.0016 0.9302 ± 0.0044 0.9846 ± 0.0026 0.9933 ± 0.0004 0.9288 ± 0.0050
AURC (↓) GCN 0.0392 ± 0.0037 0.1084 ± 0.0065 0.0442 ± 0.0020 0.0148 ± 0.0017 0.0937 ± 0.0066
GNODE 0.0439 ± 0.0034 0.1396 ± 0.0159 0.0512 ± 0.0070 0.0163 ± 0.0028 0.1154 ± 0.0235
BGCN 0.0399 ± 0.0049 0.2522 ± 0.0795 0.0704 ± 0.0102 0.0204 ± 0.0069 0.3574 ± 0.0695
ENSEMBLE 0.0543 ± 0.0038 0.1410 ± 0.0111 0.0375 ± 0.0027 0.0143 ± 0.0013 0.1007 ± 0.0076
GRAPH GP 0.1027 ± 0.0101 0.1975 ± 0.0108 OOM OOM OOM
LGNSDE (Ours) 0.0387 ± 0.0033 0.1249 ± 0.0106 0.0394 ± 0.0031 0.0148 ± 0.0013 0.1030 ± 0.0094
Accuracy (↑) GCN 0.8626 ± 0.0072 0.7720 ± 0.0051 0.8734 ± 0.0054 0.9315 ± 0.0042 0.8130 ± 0.0077
GNODE 0.8450 ± 0.0088 0.7308 ± 0.0269 0.8731 ± 0.0069 0.9247 ± 0.0041 0.7888 ± 0.0207
BGCN 0.8556 ± 0.0115 0.5530 ± 0.0992 0.8274 ± 0.0272 0.9049 ± 0.0150 0.4996 ± 0.0721
ENSEMBLE 0.8384 ± 0.0057 0.7253 ± 0.0115 0.8844 ± 0.0055 0.9280 ± 0.0039 0.7911 ± 0.0109
GRAPH GP 0.7727 ± 0.0114 0.7148 ± 0.0141 OOM OOM OOM
LGNSDE (Ours) 0.8641 ± 0.0061 0.7534 ± 0.0116 0.8830 ± 0.0058 0.9318 ± 0.0012 0.8097 ± 0.0111
Table 4: Results of the active learning experiments using a random acquisition function. The table reports AUROC, AURC, and Accuracy (Mean ± Std) for all datasets. Red indicates the best-performing model, and blue indicates the second-best-performing model.
To provide a baseline comparison with the maximum entropy acquisition function, we first evaluate the models using a random acquisition function. At each active learning round, random labels are selected to be included in the training set. Table 4 illustrates the performance of the models under this setup, showing the AUROC, AURC, and accuracy metrics across all datasets.
This setup allows us to contrast the effectiveness of random label selection with more informed selection strategies like the maximum entropy acquisition function, which we explore next.
5.4.2 Total Entropy acquisition function
Metric Model Cora Citeseer Computers Photo Pubmed
AUROC (↑) GCN 0.9831 ± 0.0013 0.9434 ± 0.0053 0.9881 ± 0.0008 0.9936 ± 0.0004 0.9248 ± 0.0129
GNODE 0.9789 ± 0.0027 0.9419 ± 0.0035 0.9838 ± 0.0017 0.9914 ± 0.0006 0.9113 ± 0.0140
BGCN 0.9830 ± 0.0026 0.8073 ± 0.0495 0.9809 ± 0.0029 0.9915 ± 0.0014 0.7169 ± 0.0529
ENSEMBLE 0.9755 ± 0.0023 0.9192 ± 0.0049 0.9867 ± 0.0005 0.9916 ± 0.0004 0.9236 ± 0.0052
GRAPH GP 0.9520 ± 0.0026 0.9122 ± 0.0030 OOM OOM OOM
LGNSDE (Ours) 0.9850 ± 0.0010 0.9426 ± 0.0057 0.9892 ± 0.0001 0.9942 ± 0.0002 0.9338 ± 0.0050
AURC (↓) GCN 0.0391 ± 0.0037 0.1104 ± 0.0065 0.0425 ± 0.0024 0.0153 ± 0.0010 0.1051 ± 0.0215
GNODE 0.0428 ± 0.0050 0.1136 ± 0.0056 0.0658 ± 0.0080 0.0206 ± 0.0022 0.1147 ± 0.0199
BGCN 0.0383 ± 0.0045 0.2849 ± 0.0595 0.0733 ± 0.0176 $0$ 0.0053 0.3716 ± 0.0681
ENSEMBLE 0.0524 ± 0.0080 0.1402 ± 0.0081 0.0353 ± 0.0016 0.0188 ± 0.0010 0.1053 ± 0.0089
GRAPH GP 0.1007 ± 0.0039 0.2022 ± 0.0043 OOM OOM OOM
LGNSDE (Ours) 0.0353 ± 0.0027 0.1199 ± 0.0110 0.0378 ± 0.0010 0.0139 ± 0.0007 0.0980 ± 0.0094
Accuracy (↑) GCN 0.8806 ± 0.0065 0.7877 ± 0.0077 0.8850 ± 0.0045 0.9355 ± 0.0020 0.7995 ± 0.0238
GNODE 0.8710 ± 0.0091 0.7836 ± 0.0086 0.8708 ± 0.0048 0.9302 ± 0.0047 0.7799 ± 0.0280
BGCN 0.8728 ± 0.0089 0.5167 ± 0.0727 0.8208 ± 0.0246 0.9104 ± 0.0199 0.4788 ± 0.0707
ENSEMBLE 0.8627 ± 0.0063 0.7530 ± 0.0075 0.8996 ± 0.0020 0.9410 ± 0.0010 0.7924 ± 0.0119
GRAPH GP 0.7868 ± 0.0061 0.7129 ± 0.0051 OOM OOM OOM
LGNSDE (Ours) 0.8889 ± 0.0067 0.7826 ± 0.0099 0.8984 ± 0.0021 0.9381 ± 0.0017 0.8208 ± 0.0111
Table 5: Results of the active learning experiments using a maximum entropy acquisition function. The table reports AUROC, AURC, and Accuracy (Mean ± Std) for all datasets. Red indicates the best-performing model, and blue indicates the second-best-performing model.
Now, we evaluate the models using a total entropy acquisition function, where labels with the highest uncertainty are selected at each active learning round. Table 5 shows the AUROC, AURC, and accuracy metrics for all datasets. Our model, LGNSDE, performs significantly better with the total entropy acquisition function compared to the random acquisition function (Table 4). It achieves the highest AUROC on Cora, Computers, Photo, and Pubmed, showing that it effectively selects the most informative points when uncertainty is used as a guide. In contrast, the random function led to more mixed results across these datasets. For the AURC metric, where lower is better, LGNSDE shows clear improvements with the entropy strategy, particularly on Cora, Computers, and Pubmed. This suggests that our model is better at reducing classification errors when it focuses on uncertain points, compared to the random selection method. In terms of accuracy, LGNSDE also sees gains under the total entropy function. For example, accuracy on Cora improves from 0.8641 to 0.8889, and on Pubmed from 0.8097 to 0.8208, compared to the random acquisition setup. This shows that using uncertainty to guide label selection leads to better overall performance.
In summary, the total entropy acquisition function helps LGNSDE perform more effectively, particularly by selecting more informative data points than random selection, resulting in higher accuracy and better uncertainty management. For details on the effects of the hyperparameters and additional benchmark experiments, please refer to Appendix C.
6 Related Work
Uncertainty quantification in GNNs has recently gained attention, with contributions from Bayesian GNNs (Hasanzadeh et al., 2020), ensemble-based methods (Lin et al., 2022), and Gaussian Processes on graphs (Borovitskiy et al., 2021; Sáez de Ocáriz Borde et al., 2024). We benchmark our LGNSDE framework against these methods and show improved performance across tasks, demonstrating more flexibility in capturing both aleatoric and epistemic uncertainty through a dynamic stochastic framework.
The work of Bishnoi et al. (2023) introduces stochastic elements into graph learning but is restricted to learning scalar parameters for an SDE, unlike our method, which models the graph evolution directly as an SDE. Similarly, Poli et al. (2021) attempt to propose Graph Neural SDEs with a different formulation that uses a finite-dimensional KL divergence in the ELBO instead of an infinite-dimensional version. They also use prior distributions instead of prior processes, which are better suited for modelling continuous dynamical systems. Hence, constructing a different method. Moreover, they do not provide theoretical analysis or thoroughly explore the model’s uncertainty estimation capabilities beyond limited toy experiments.
The framework proposed by Lin et al. (2024) employs stochastic partial differential equations (SPDEs) to model message passing for uncertainty estimation, leveraging a novel Q-Wiener process to propagate uncertainty directly within the graph diffusion process. While their approach emphasizes diffusion-based uncertainty, our method adopts a Bayesian framework with SDEs, focusing on quantifying uncertainty in the latent space instead. Furthermore, the work by Stadler et al. (2021) on Graph Posterior Networks (GPN) takes a Bayesian approach to uncertainty estimation, focusing on interdependent nodes in graph-structured data. Their model explicitly performs posterior updates for node-level classification by leveraging Dirichlet distributions and performs well in uncertainty-sensitive tasks. However, GPNs do not capture the dynamic evolution of node embeddings over time as our LGNSDE does. Lastly, (Zhao et al., 2020) work on uncertainty-aware semi-supervised learning for graphs introduces a method which uses belief theory to quantify uncertainty types like vacuity and dissonance. Their focus on semi-supervised learning and belief theory-based uncertainty contrasts with our SDE-based approach, where uncertainty is captured through variance in latent representations and modelled using Bayesian updates.
@@ 7 Conclusions and Future Work
We have introduced Latent Graph Neural Stochastic Differential Equations (LGNSDE), a novel framework designed to quantify uncertainty in graph-structured data. By leveraging both epistemic and aleatoric uncertainty through a Bayesian prior-posterior mechanism and Brownian motion, LGNSDE provides meaningful uncertainty estimates. Theoretical guarantees demonstrate that our model ensures well-posedness, variance bounds, and robustness to small perturbations in inputs. Empirically, LGNSDE performs competitively across a range of benchmarks, excelling in out-of-distribution detection, noise robustness, and active learning tasks. Future directions include exploring higher-order SDEs, optimizing computational efficiency, and applying the model to real-world systems requiring reliable uncertainty quantification such as recommendation systems, drug discovery or dynamic networks.
@@ A.1 Notation
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ denote a graph with node set $\mathcal{V}$ and edge set $\mathcal{E}$. The node feature matrix at time $t$ is $\mathbf{H}(t)\in\mathbb{R}^{n\times d}$, where $n$ is the number of nodes and $d$ is the feature dimension. The evolution of $\mathbf{H}(t)$ is described by a Graph Neural Stochastic Differential Equation, with drift function $\mathbf{F}_{\mathcal{G}}(\mathbf{H}(t),t,\bm{\theta})$ and diffusion function $\mathbf{G}_{\mathcal{G}}(\mathbf{H}(t),t)$. Here, $\mathbf{F}_{\mathcal{G}}$ depends on the graph $\mathcal{G}$, the node features $\mathbf{H}(t)$, time $t$, and parameters $\bm{\theta}$. The diffusion function $\mathbf{G}_{\mathcal{G}}$ depends on $\mathcal{G}$ and $\mathbf{H}(t)$ but not on $\bm{\theta}$, as in practice, this is usually a constant function. The randomness is introduced through the Brownian motion $\mathbf{W}(t)$.
The constants $L_{f}$ and $L_{g}$ are Lipschitz constants for the drift and diffusion functions, respectively, ensuring the existence and uniqueness of the solution to the GNSDE. The linear growth condition is controlled by a constant $K$, preventing unbounded growth in $\mathbf{F}_{\mathcal{G}}$ and $\mathbf{G}_{\mathcal{G}}$. Finally, $\text{Var}(\mathbf{H}(t))$ represents the variance of the node features, capturing the aleatoric uncertainty in the system, which is also reflected in the variance of the model output $\mathbf{\hat{y}}(t)=\mathbf{h}(\mathbf{H}(t))$.
@@ A.2 Technical Assumptions
$\|\mathbf{F}_{\mathcal{G}}(\mathbf{H}_{1}(t),t,\bm{\theta})-\mathbf{F}_{\mathcal{G}}(\mathbf{H}_{2}(t),t,\bm{\theta})\|_{F}$ $\leq L_{f}\|\mathbf{H}_{1}(t)-\mathbf{H}_{2}(t)\|_{F}$ (3)
@@
$\|\mathbf{F}_{\mathcal{G}}(\mathbf{H}(t),t,\bm{\theta})\|_{F}^{2}+\|\mathbf{G}_{\mathcal{G}}(\mathbf{H}(t),t)\|_{F}^{2}\leq K(1+\|\mathbf{H}(t)\|_{F}^{2}),$
@@ A.3 Proofs
See 1
@@
Using the result in 1 (Lin et al. (2024)), it follows that the Lipschitz conditions of $\mathbf{F}_{\mathcal{G}}$ and $\mathbf{G}_{\mathcal{G}}$ ensure the existence and uniqueness of a mild solution $\mathbf{H}(t)$ to the GNSDE.
@@
Finally, assuming that the model output $\mathbf{\hat{y}}(t)$ is a function of the latent state $\mathbf{H}(t)$, $\mathbf{\hat{y}}(t)=\mathbf{h}(\mathbf{H}(t))$, where $\mathbf{h}:\mathbb{R}^{n\times d}\rightarrow\mathbb{R}^{n\times p}$ is a smooth function, we can apply Itô’s Lemma as follows:
$dy(t)=h^{\prime}(\mathbf{H}(t))\left[\mathbf{F}_{\mathcal{G}}(\mathbf{H}(t),t,\bm{\theta})\,dt+\mathbf{G}_{\mathcal{G}}(\mathbf{H}(t),t)\,d\mathbf{W}(t)\right]+\frac{1}{2}h^{\prime\prime}(\mathbf{H}(t))\mathbf{G}_{\mathcal{G}}(\mathbf{H}(t),t)^{2}\,dt.$
For the variance of $\mathbf{\hat{y}}(t)$, we focus on the term involving $\mathbf{G}_{\mathcal{G}}(\mathbf{H}(t),t)\,d\mathbf{W}(t)$:
@@
See 2
@@
$d(\delta\mathbf{H}(t))=\left[\mathbf{F}_{\mathcal{G}}(\mathbf{H}_{1}(t),t,\bm{\theta})-\mathbf{F}_{\mathcal{G}}(\mathbf{H}_{2}(t),t,\bm{\theta})\right]dt+\left[\mathbf{G}_{\mathcal{G}}(\mathbf{H}_{1}(t),t)-\mathbf{G}_{\mathcal{G}}(\mathbf{H}_{2}(t),t)\right]d\mathbf{W}(t).$
@@
$d(\text{tr}(\delta\mathbf{H}(t)^{\top}\delta\mathbf{H}(t)))$ $=2\text{tr}\left(\delta\mathbf{H}(t)^{\top}\left[\mathbf{F}_{\mathcal{G}}(\mathbf{H}_{1}(t),t,\bm{\theta})-\mathbf{F}_{\mathcal{G}}(\mathbf{H}_{2}(t),t,\bm{\theta})\right]\right)dt$
@@
$\mathbb{E}[d(\text{tr}(\delta\mathbf{H}(t)^{\top}\delta\mathbf{H}(t)))]$ $=\mathbb{E}\left[2\text{tr}(\delta\mathbf{H}(t)^{\top}[\mathbf{F}_{\mathcal{G}}(\mathbf{H}_{1}(t),t,\bm{\theta})-\mathbf{F}_{\mathcal{G}}(\mathbf{H}_{2}(t),t,\bm{\theta})])\right]\,dt$
@@
A.4 LGNSDE as a Continuous Representation of Graph ResNet with Stochastic Noise Insertion
Consider a Latent Graph Neural Stochastic Differential Equation (LGNSDE) represented as